OpenStudy (anonymous):

Evaluate expressions. Give exact values. Sec(squared)135 degrees + 2sin210 degrees

OpenStudy (anonymous):

cos 45 = 1/sqrt2 , cos 135 = -1/sqrt 2 so sec 135 = -sqrt2 and squared = 2 sin210 = -sin 30 = -1/2, and 2sin210 = -1 answer = 2 -1 = 1